The teenager company makes and sales skate boards at an average price of $70 each.During the past year they sold 4000 of these skateboards. The company believes that the prices that the price elasticity for this product is about -2.5. If its price decrease to $ 63.what should be the quantity sold?will revenue Increase ?why?
1
Expert's answer
2018-10-24T11:12:09-0400

price elasticity formula is: E_p^Q = ΔQ/Q : ΔP/P , when E_p^Q – elasticity in price, ΔQ – change of quantity (Q next year – Q past year), ΔP - change of price (P next year – P past year )

if we substitute the available values we take:

(Qn-4000)/4000 : (63-70)/70 = -2.5


(Qn-4000)/4000 * ((-70)/7) = -2.5

(-10(Qn-4000))/4000 = - 2.5

(Qn-4000) = 1000

Qn = 5000


Answer: Quantity sold is 5000 skateboards. As we can see, revenue is increase after price decrease. This can be explained by the law of demand. When the price of a product decreases, the demand for it increases, which ultimately leads to an increase in sales revenue due to an increase in the number of products (goods) sold.

At the same time, we note that the skateboard product itself is highly elastic due to the fact that it is not critical for the consumer (the elasticity of this product is a negative value).